### v1

# Large Language Models for Design Structure Matrix Optimization

Abstract
In complex engineering systems, the interdependencies among components or development activities are often modeled and analyzed using Design Structure Matrix (DSM). Reorganizing elements within a DSM to minimize feedback loops and enhance modularity or process efficiency constitutes a challenging combinatorial optimization (CO) problem in engineering design and operations. As problem sizes increase and dependency networks become more intricate, traditional optimization methods that solely use mathematical heuristics often fail to capture the contextual nuances and struggle to deliver effective solutions. In this study, we explore the potential of Large Language Models (LLMs) for helping solve such CO problems by leveraging their capabilities for advanced reasoning and contextual understanding. We propose a novel LLM-based framework that integrates network topology with contextual domain knowledge for iterative optimization of DSM element sequencing - a common CO problem. Experiments on various DSM cases show that our method consistently achieves faster convergence and superior solution quality compared to both stochastic and deterministic baselines. Notably, we find that incorporating contextual domain knowledge significantly enhances optimization performance regardless of the chosen LLM backbone. These findings highlight the potential of LLMs to solve complex engineering CO problems by combining semantic and mathematical reasoning. This approach paves the way towards a new paradigm in LLM-based engineering design optimization.
Keywords Large Language Models  $\cdot$ Artificial Intelligence  $\cdot$ Design Structure Matrix  $\cdot$ Combinatorial Optimization  $\cdot$ Engineering Design  $\cdot$ Systems Engineering
1 Introduction
Design Structure Matrix (DSM) is a well-established modeling tool in engineering design [1, 2]. It encodes relationships among elements, such as activities, components, or parameters, in a compact matrix format [2]. Reordering the sequence of nodes in a DSM to minimize feedback loops and improve modularity can improve system performance, reduce development risk, and increase design efficiency [1, 3]. Figure 1 illustrates an example of a design activity DSM before and after sequencing [4]. DSM sequencing is a typical combinatorial optimization (CO) problem, and also an NP-hard problem. As matrix sizes and dependency intricacy increase, traditional optimization methods that rely on heuristics and rule-based algorithms often struggle to find effective solutions [5, 6].
Figure 1: Illustration of a Design Activity DSM: (A) Pre-Sequencing; (B) Post-Sequencing
Recent advancements in Large Language Models (LLMs) have demonstrated their powerful capabilities in natural language generation, semantic understanding, instruction following, and complex reasoning [7, 8]. Furthermore, studies have shown that LLMs can be used for continuous and concrete optimization [9, 10, 11, 12]. For instance, researchers from DeepMind utilized LLMs as optimizers and evaluated their effectiveness on classic CO problems, such as Traveling Salesman Problems [12]. In the field of engineering, researchers are also exploring generative artificial intelligence (GenAI) to address engineering optimization problems [13]. In addition, previous research has shown that LLMs possess extensive domain knowledge learned during pretraining on diverse engineering and scientific data resources, which enhances their applicability in engineering contexts [14, 15, 16, 17].
Therefore, the ability of LLMs to combine mathematical and semantic reasoning, along with their possession of extensive domain knowledge, motivated us to explore their potential for DSM optimization tasks in engineering design, where contextual information is vital for identifying high-quality solutions. Our hypotheses are: (1) LLMs can be effectively applied to DSM-based combinatorial optimization tasks, and (2) incorporating contextual domain knowledge can further enhance LLM performance by supporting mathematical reasoning with semantic insights.
This paradigm, which leverages both semantic and mathematical reasoning, introduces a novel approach to combinatorial optimization beyond the reach of traditional purely mathematical methods. To validate the proposed method, we conduct experiments on various engineering-related DSM cases collected from peer-reviewed literature. The experimental results demonstrate the practical applicability of our proposed LLM-based approach for combinatorial optimization of DSM. The main contributions of this paper are summarized as follows:
  1. (1)
    We propose a novel LLM-based optimization framework that integrates network topology and contextual domain knowledge for DSM sequencing optimization.
  2. (2)
    We perform a comprehensive benchmarking analysis against both stochastic and deterministic algorithms, and systematically evaluate the impact of different backbone LLMs, validating the robustness and adaptability of our method across various engineering DSM cases.
  3. (3)
    We show that incorporating contextual domain knowledge significantly improves both solution quality and convergence speed in our method. In addition, to the best of our knowledge, this is the first study to integrate LLMs and DSM for engineering optimization.
This paper is organized as follows. Section 2 reviews related work on DSM and LLMs for optimization. Section 3 presents our proposed LLM-based framework in detail. Section 4 describes the data, experiments, and benchmark methods. Section 5 presents the experimental results and discusses key insights and practical implications for engineering design. Section 6 summarizes the limitations of the study and discusses potential directions for future work. Finally, Section 7 concludes the paper.
2 Background and Related Work
2.1 Design Structure Matrix for Engineering Design
The DSM is an essential tool that models the interactions and dependencies among elements within complex engineering systems [1]. Originating from Donald Steward’s pioneering work [2], DSMs provide a systematic framework to represent and manage dependencies among elements in complex engineering systems, using a compact matrix format. Each row and column in the matrix correspond to a specific component, activity, or parameter, while the matrix entries illustrate the state and strength of their interdependencies. DSM methodologies have evolved significantly, with researchers proposing advanced methods for managing complex engineering design processes, improving modularity, and enhancing overall system efficiency [5]. Since then, DSMs have seen significant development in both academia and industry. DSMs have been successfully applied across many disciplines for engineering design, such as automotive [18, 19], semiconductor [20], aerospace [21], factory equipment [22], and more [23, 24, 25, 26, 27].
DSM can be categorized primarily into four types: component-based, team-based, activity-based, and parameter-based [28]. The first two types are static DSMs, whereas the other two are time-based DSMs. Each category addresses specific aspects of engineering design, ranging from structuring task flows and team interactions to the relationships between technical parameters of system components. Static DSMs are typically analyzed using clustering algorithms, which identify modular groupings to minimize interdependencies and facilitate parallel development. Time-based DSMs are usually analyzed using sequencing algorithms to streamline processes, reduce iterative feedback loops, and optimize the logical progression of activities.
The optimization of DSMs in engineering design are typical CO problems, requiring efficient and effective methods to obtain high-quality solutions. Researchers have explored various algorithms for DSM applications, including heuristic, exact and hybrid approaches [6, 29, 30, 31, 32, 33]. These previous works have continuously advanced mathematical methods to enhance the performance and scalability of DSM optimization. However, while DSMs abstractly represent real-world engineering problems, they often encode rich contextual domain knowledge that can be valuable for guiding optimization. With the capabilities of large-scale foundational AI models, such information can be effectively incorporated into the optimization process. Thus, this paper contributes to this field by leveraging LLMs for DSM optimization, considering both semantic and mathematical reasoning.
2.2 Large Language Models for Combinatorial Optimization
Combinatorial optimization problems are frequently encountered across fields, where finding an optimal solution from a finite set often drives improvements in efficiency, cost, and performance [34]. For instance, applications such as DNA barcoding and DNA assembly in synthetic biology [35], as well as job scheduling in manufacturing [36], rely heavily on effective solutions of CO problems. However, due to their NP-hard nature, these problems present substantial challenges, especially as complexity increases with larger problem sizes and more intricate dependencies. Traditionally, real-world CO problems are usually approached through the following process: the problem is first modeled formally and mathematically, then solved using specific optimization algorithms or heuristics, and finally interpreted within the context of practical engineering [37]. This separation between problem-solving and interpretation stages limits the ability to capture the contextual nuances that are often essential for optimizing real-world problems.
Recent advancements have significantly enhanced the capabilities of LLMs [7, 8]. LLMs acquire extensive common-sense and domain-specific knowledge during their pretraining on large-scale datasets, including publicly available code repositories, natural language corpora, and other multimodal data. This broad foundation is further refined through fine-tuning on various downstream tasks, enabling LLMs to adapt to specific applications. State-of-the-art models such as DeepSeek, LLaMA, Claude, and GPT-series models have demonstrated remarkable performance in understanding natural language, complex reasoning, and generalizing across a wide range of tasks [38, 39, 40, 41].
LLMs have recently been leveraged to address CO problems [9, 10, 11, 12]. Early explorations utilized prompt engineering techniques to guide LLMs in solving classical CO problems, such as the traveling salesman problem and job-shop scheduling, demonstrating the feasibility of using LLMs for direct solution generation [10, 12]. Later, inspired by automatic heuristic generation, researchers proposed hybrid approaches that combine LLMs with evolutionary computation to generate, adapt, and refine optimization heuristics [9, 11]. Compared with traditional optimization algorithms, these approaches benefit from LLMs’ flexible reasoning and knowledge-driven heuristic generation, enabling generalization across problem variants without handcrafted solvers.
However, most of these studies simplify CO problems into a formal representation, overlooking contextual domain-specific knowledge that is critical in real-world applications. A recent study [42] demonstrated that LLMs can generate interpretable DSM structures directly from textual descriptions of engineering processes. Such exploratory research highlights that LLMs’ ability to capture relevant domain knowledge [14, 15, 43]. Therefore, the ability of LLMs to integrate both mathematical and semantic reasoning, along with their access to extensive pretrained knowledge, motivates us to explore their potential for solving CO problems in engineering design.
3 Methodology
In this section, we present our proposed LLM-based approach in detail and apply it to solve DSM element sequencing problems, aiming to reduce backward dependencies and enhance system modularity in engineering workflows. Overall, the approach is designed to harness the generative and reasoning capabilities of LLMs, combining them with domain-specific contextual knowledge and a formal evaluation mechanism to iteratively guide solution optimization.
The process begins with the initialization of a solution that is randomly sampled from the total solution space. Each solution is then evaluated using a predefined evaluator that quantifies its quality based on specific criteria. Guided by this evaluation, the approach iteratively updates the solution base by leveraging the in-context learning capabilities of LLMs. Prompts are crafted using both the network’s mathematical representation and contextual domain knowledge in natural language descriptions, along with meta-instructions and selected historical solutions. In each iteration, the LLM generates a new candidate solution, which is then validated and evaluated. The solution, along with its evaluation results, is appended to the solution base. Once the termination condition is met, the approach outputs the best-performing solution as the final result. An overview of the proposed approach is shown in Figure 2.
Figure 2: Overview of the Proposed LLM-based Approach
3.1 Initialization and Solution Sampling
The solution base is a core component of the approach, serving three primary purposes: (1) storing all previously explored solutions along with their evaluation results, (2) providing historical solutions for few-shot learning to guide the LLM’s generation in subsequent iterations, and (3) returning the top-performing solution as the final output at the end of the optimization process.
The initialization phase begins by randomly sampling a valid solution from the entire solution space. In the context of DSM sequencing, each solution corresponds to a complete and non-repetitive permutation of nodes (see Figure 2). This initial solution is evaluated using the predefined evaluator and added to the solution base. During subsequent iterations, a sampling strategy is employed to construct a candidate solution set. Specifically, we select the top $K_{p}$ highest-scoring solutions and randomly sample additional $K_{q}$ solutions from the remaining $K_{n}-K_{p}$ entries in the solution base, where $K_{n}$ denotes the total number of stored solutions. The combined set of $K_{p}+K_{q}$ solutions is then used to construct prompts that guide the next round of LLM-based generation. The values of $K_{p}$ and $K_{q}$ are tunable hyperparameters that balance exploitation and exploration in the search process.
3.2 LLM-driven Generative Optimization
In each iteration of the optimization process, the backend LLM is prompted with a structured input that incorporates both mathematical representations of the DSM and contextual knowledge related to the given DSM. The prompt includes the following four key components:
(i) Typological Information: This provides a mathematical description of the DSM. While multiple representations can be used to describe a complex network (e.g., an adjacency matrix, a dependency list, or an edge list), we adopt the edge list format for its simplicity and compatibility with sequence-based reasoning. To mitigate any ordering bias, all edges are randomly shuffled before being included in the prompt.
(ii) Contextual Domain Knowledge: This component includes the semantic information associated with each node (e.g., activity names, component functions) and an overall textual description of the network. These natural language elements convey the engineering context behind the DSM structure, enabling the LLM to reason beyond abstract graph features. For example, in an activity-based DSM, nodes represent distinct phases of the engineering design process.
(iii) Meta-instructions: To guide the generation behavior of the LLM, we adopt several frequently used prompt engineering strategies [44], including role-playing, task specification, and output format specification. These instructions help structure the LLM’s responses and ensure syntactic and semantic consistency.
(iv) Selected Historical Solutions: As described in the previous section, a curated set of $K_{p}+K_{q}$ solutions are then sampled from the solution base and presented to the LLM for few-shot learning. This allows the model to generalize from prior high-quality and diverse solutions.
Upon receiving the complete input, the LLM generates a new candidate solution, which is then validated by a checker to ensure that it includes all nodes exactly once and adheres to the required sequencing constraints. Valid solutions are then evaluated using the predefined scoring function and subsequently appended to the solution base for use in future iterations. Details of the input information are provided in Appendix 1.
3.3 Evaluation of Generated Solutions
To evaluate solution quality, each new sequence generated by the LLM is assessed upon creation. In DSM sequencing, the objective is to reorder the rows and columns of the DSM to minimize the number of feedback loops. The resulting score is then used to rank solutions in the solution base and to guide sampling and prompt construction in subsequent iterations. For optimization problems other than DSM sequencing, the evaluation criteria can be adapted based on the structural characteristics and design objectives of the specific problem. For instance, in modularity optimization (i.e., clustering), the metric may focus on maximizing intra-module cohesion while minimizing inter-module dependencies.
When the predefined number of iterations is reached or the evaluation score meets a known optimal threshold, the iteration process is terminated. At that point, the best-performing solution identified is returned as the final output.
4 Experiments
4.1 Data
We curated four DSM cases for experiments, which are categorized into two types following established classifications [28]: (1) Activity-based DSMs, which represent the input-output relationships among tasks or activities within a project; and (2) Parameter-based DSMs, which illustrate the precedence and interdependencies among design parameters.
For activity-based DSMs, we collected two cases:
  • The Unmanned Combat Aerial Vehicle (UCAV) DSM includes 12 conceptual design activities conducted at Boeing [45].
  • The Microfilm Cartridge DSM was derived from Kodak’s Cheetah project and consists of 13 major development tasks [46].
Both DSMs were constructed through structured interviews with domain experts and were subsequently reviewed and validated to ensure accuracy and consistency.
For parameter-based DSMs, we collected two cases:
  • The Heat Exchanger DSM, which includes 17 components related to core thermal exchange functions [4].
  • The Automobile Brake System DSM, which comprises 14 design parameters reflecting braking mechanisms and their interdependencies [47].
Both DSMs were developed by first identifying key product components, followed by interviews with the respective design engineers to define parameters and establish directional relationships.
From the original references, we extracted three types of information for each DSM: (1) the names of all nodes (activities or parameters), (2) the edge list derived from the adjacency matrix, and (3) a natural language description summarizing the system-level context of each network. The data were maintained in their original structure to ensure fidelity and reproducibility. Data formatting details are provided in Appendix 1.
The characteristics of the four DSMs are summarized in Table 1. Overall, the number of nodes $N$ ranges from 12 to 17, and the number of edges $E$ varies between 32 and 47. We also report network-level metrics including diameter, density, average degree, clustering coefficient, and average path length to present topological complexity. For example, the UCAV DSM exhibits a high network density of 0.712 and a clustering coefficient of 0.773, reflecting tightly coupled design tasks. In contrast, the Heat Exchanger DSM is more sparsely connected, with clearer modular separations among design parameters.
Table 1: Characteristics of Four DSMs
N E Network Diameter Network Density Average Degree Clustering Coefficient Average Path Length
Activity-Based DSMs Unmanned Aerial Vehicle [45] Microfilm Cartridge [46] 12 47 2 0.712 7.833 0.773 1.288
13 41 3 0.526 6.308 0.682 1.577
Parameter-Based DSMs Heat Exchanger [4] Automobile Brake System [47] 17 41 7 0.302 4.824 0.457 2.397
14 32 4 0.352 4.571 0.414 1.824
4.2 Evaluation Metrics
To evaluate the quality and effectiveness of solutions generated during the optimization process, we employ the following metrics aligned with the goals of DSM sequencing tasks. The choice of metrics is consistent with prior studies [1]. Throughout the iterative optimization process, we use the number of feedback loops (i.e., non-zero entries above the diagonal in the adjacency matrix) as the primary metric to guide solution selection and LLM-driven generation.
Given a sequence, we first construct an $n\times n$ asymmetric adjacency matrix $A$ based on the edge list of the network, where $n$ denotes the number of nodes. Each binary entry $a_{ij}$ in the matrix indicates whether node $i$ depends on node $j$, with $a_{ij}=1$ representing a directed edge from $j$ to $i$. The sequencing task involves reordering the nodes such that the number of non-zero entries above the main diagonal in $A$ is minimized. These entries correspond to feedback dependencies that disrupt ideal information or task flow. Formally, given a sequence $s$, the evaluation score is computed as:
Score(s) = ∑_i=1^n ∑_j=i+1^n a_ij
Lower values correspond to improved modularization or reduced rework potential in engineering workflows. Given the heuristic and iterative nature of our approach, the number of feedback loops is computed at each iteration for every valid solution. This allows us to track convergence and identify high-quality solutions progressively. For a fair comparison across methods, we also report the best achieved value within a fixed number of trials for each configuration in our experiments.
During performance benchmarking, we complement this primary metric with statistical reporting of mean and standard deviation over multiple trials. For convergence analysis, we record the number of unique sequences explored before reaching the optimal or near-optimal score, thereby capturing both efficiency and solution quality.
4.3 Experiment Setup
To ensure the robustness of results, each method was executed 10 times with different random seeds. In the solution sampling process, we set the top-performing selection parameter $K_{p}=5$ and the random sampling parameter $K_{q}=5$. Each node in the DSM was assigned a unique identifier, composed of five randomly generated alphanumeric characters, to ensure independence from domain-specific naming. The maximum number of iterations for all experiments was fixed at 20. Claude-3.5-Sonnet-20241022 was used as the main backbone LLM, and all model parameters were kept at their default settings [38].
4.4 Baseline Methods
To benchmark the performance of our proposed approach, we compare it against both stochastic and deterministic optimization approaches commonly used for CO problems and DSM sequencing tasks.
Stochastic Methods. We adopt the classic Genetic Algorithm (GA), a widely used probabilistic optimization method [48], implemented using the DEAP library [49]. Three configurations were tested to explore different optimization strategies:
  1. (1)
    Exploration-focused: Prioritizes diversity to avoid premature convergence.
  2. (2)
    Exploitation-focused: Focuses on intensifying search around high-performing regions.
  3. (3)
    Balanced setting: Seeks a compromise between exploration and exploitation.
Detailed settings are provided in Appendix 2.
Deterministic Methods. Deterministic baselines reorder nodes based on fixed structural metrics. Since the calculation of each measure is deterministic, the resulting solution is consistent. Five representative deterministic algorithms were selected:
  1. (1)
    Out-In Degree [50]: Orders nodes by the difference between out-degree and in-degree.
  2. (2)
    Eigenvector [51]: Uses the Perron vector of the adjacency matrix to rank nodes.
  3. (3)
    Walk-based (Exponential) [52]:Computes $F(\mathbf{A})=\exp(\mathbf{A})$, capturing long-range connectivity.
  4. (4)
    Walk-based (Resolvent) [53]: Computes $F(\mathbf{A})=(\mathbf{I}-\delta\mathbf{A})^{-1}$, where $\delta=0.025$ represents the probability that a message will successfully traverse an edge. The value of $\delta$ is adopted based on prior work.
  5. (5)
    Visibility [54]: Computes $F(\mathbf{A})=\sum_{k=0}^{n}\mathbf{A}^{k}$, aggregating all path-based dependencies, then binarizes the result.
For deterministic methods (3), (4), and (5), nodes are ranked based on the row sums of F(A); ties are resolved using column sums.
Ablation Setting. To assess the role of contextual domain knowledge, we implement a variant of our method that removes all semantic information from the input prompt. In this version, the LLM is guided solely by the DSM’s topological structure, without node names or network descriptions. This setup enables us to isolate the impact of domain knowledge on optimization performance.
5 Results and Discussion
5.1 Performance Comparison with Benchmarks
Using the number of feedback loops as the metric (described in Section 4.2), we compared the performance of our proposed LLM-based method against both stochastic and deterministic baselines across four DSM cases, as shown in Table 2. The results demonstrate the superiority of our approach in terms of solution quality, particularly when contextual domain knowledge is incorporated.
Our methods were evaluated in three configurations: single-trial, 5-trial, and 20-trial runs, each tested with and without the inclusion of contextual domain knowledge. Among deterministic baselines, which are inherently single-trial and produce consistent outputs, our single-trial method with knowledge consistently outperformed all competitors, with the sole exception (Visibility-based method), which matched our performance in the Automobile Brake System case. In that same case, our method achieved identical best results when run under the 5-trial and 20-trial configurations.
In comparison with stochastic algorithms, our 20-trial LLM-based method consistently outperformed all GA variants across all DSM cases. It achieved the optimal solution in three out of four cases. In the remaining case (Heat Exchanger DSM), our method produced a solution that closely approached the optimum. Notably, these results were obtained within a relatively limited iteration budget of only 20 trials. Even in the 5-trial setting, our method outperformed all three GA configurations in three cases. These results validate the effectiveness of the proposed approach in producing high-quality solutions with relatively few attempts.
Table 2: Performance Comparison with Baseline Methods
Methods Activity-Based DSMs Parameter-Based DSMs
Unmanned Aerial Vehicle Microfilm Cartridge Heat Exchanger Automobile Brake System
Stochastic Methods [^1] Consistent with Figure 3, all three GA settings show optimization performance with number of unique sequences explored = 10,000.
GA (Exploration-focused setting) 6.0±0.0 8.1±0.3 5.7±1.0 3.8±0.7
GA (Exploitation-focused setting) 7.4±1.5 9.9±1.4 7.6±1.4 6.6±1.7
GA (Balanced setting) 6.0±0.0 8.4±0.5 6.2±1.2 4.2±1.2
Deterministic Methods [^2] All deterministic methods can be regarded as single-trial approaches, as they produce consistent outcomes for identical inputs across repeated executions. However, certain nodes might share the same measures. For these nodes, a random order is applied, and the final statistical evaluation results are obtained from 10 runs.
Out-In Degree [50] 10.0±0.0 12.3±0.5 10.3±0.6 5.7±0.9
Eigenvector [51] 15.0±0.0 13.8±0.7 13.1±1.7 11.1±1.4
Walk-based (Exponential) [52] 15.0±0.0 12.0±0.0 8.0±0.0 11.0±0.0
Walk-based (Resolvent) [53] 9.0±0.0 12.0±0.0 8.0±0.0 11.0±0.0
Visibility [54] 25.6±2.2 8.8±0.7 6.0±1.3 3.0±0.0
LLM-driven Methods (Ours)
Single-trial with knowledge 6.6±0.7 8.0±0.0 4.8±0.6 4.4±0.9
Single-trial without knowledge 11.9±3.4 8.1±0.3 5.8±1.1 5.9±1.4
5-trial with knowledge 6.1±0.3 8.0±0.0 4.0±0.4 3.0±0.0
5-trial without knowledge 7.5±0.9 8.0±0.0 4.9±0.5 4.0±1.0
20-trial with knowledge 6.0±0.0 8.0±0.0 3.6±0.5 3.0±0.0
20-trial without knowledge 6.4±0.7 8.0±0.0 4.1±0.3 3.4±0.7
5.2 Convergence Analysis with Stochastic Algorithms
We evaluated the convergence behavior of our LLM-based methods in comparison with three configurations of GA, as shown in Figure 3. Due to their differing operational mechanisms, each GA iteration explores multiple unique solutions based on population dynamics and stochastic operations. In contrast, our LLM-based approach generates one unique solution per iteration. To ensure a fair comparison, we standardized convergence speed in terms of the number of unique solutions explored, rather than iteration count. Although the maximum generation number for GA was set to 2,000 (refer to Appendix 2), the comparison in Figure 3 focuses on the first 10,000 unique solutions to improve clarity of visualization.
Across all four DSM cases, our methods consistently demonstrate superior convergence speed relative to all GA variants. The LLM-based methods reach significantly lower objective values with far fewer explored sequences, highlighting their efficiency. This convergence advantage is evident across all cases. In each scenario, our method identifies a relatively high-quality initial solution and then continues to refine it through iterative improvement. This supports the hypothesis that LLMs can perform efficient in-context learning, adapting and optimizing based on previously observed sequences.
Moreover, the results further highlight the value of contextual domain knowledge in improving the convergence behavior. In Figures 3A (UCAV), 3C (Heat Exchanger), and 3D (Brake System), the version of our method that incorporates domain knowledge (red line) consistently outperforms the variant without such knowledge (gray line), both in terms of convergence speed and final solution quality. This indicates that LLMs are capable of effectively combining semantic reasoning with structural optimization to solve CO problems more intelligently. An interesting exception appears in Figure 3B (Microfilm Cartridge), where both variants of the LLM-based method reach the optimal solution in the first step. As a result, the two lines overlap entirely, reflecting identical performance in this case. This suggests that, for some DSMs with relatively less complexity or clearer structure, even minimal attempts may suffice for the LLM to infer optimal solutions.
Figure 3: Convergence Analysis
5.3 Ablation Study on the Backbone LLM
To evaluate the generalizability and adaptability of our approach across different backbone LLMs, we conducted an ablation study. In addition to our main model, Claude-3.5-Sonnet-20241022 [38], we tested four popular models, including Mixtral-7x8B [55], Llama3-70B [40], DeepSeek-V3-20250324 [39], and GPT-4-Turbo [41]. We evaluated each backbone under three settings: single-trial, 5-trial, and 20-trial runs, and we also examined performance both with and without contextual domain knowledge.
As shown in Table 3, Claude-3.5-Sonnet consistently achieved the best overall performance across all four DSM cases and all trial settings, particularly when domain knowledge was included. Its average scores were not only the lowest but also the most stable across activity-based and parameter-based DSMs. For instance, in the 20-trial setting, Claude-3.5-Sonnet achieved optimal results on three cases with no variance. When using open-source LLMs as the backbone model, however, the best-performing model varied across different cases. Specifically, DeepSeek-V3 performed best on the Microfilm Cartridge and Brake System cases, while Llama3-70B outperformed DeepSeek-V3 on the Heat Exchanger and UCAV cases. These results suggest that when adopting open-source LLMs, it may be beneficial to involve the routing mechanism that dynamically selects or ensembles models based on the characteristics of the problem instance to overcome the constraints of individual LLMs [56]. The open-source LLMs can serve as effective and practical alternatives, especially when computational budget or deployment constraints should be considered. While we recommend Claude-3.5-Sonnet as the default choice for optimal performance, open-source models such as DeepSeek-V3 or Llama3-70B may offer a viable balance between effectiveness, flexibility, and accessibility.
The results also indicate the critical role of contextual domain knowledge. Across all LLMs, solutions generated with domain knowledge consistently outperformed their knowledge-ablated versions. For example, DeepSeek-V3 achieved 3.1±0.3 with knowledge on the Brake System DSM under 20 trials, compared to 4.0±1.3 without knowledge. This supports our earlier analysis, highlighting that incorporating contextual semantic reasoning significantly enhances solution quality and convergence performance in DSM optimization.
Table 3: Ablation on the Backbone LLM
# of Trial Backbone LLM Knowledge Activity-Based DSMs Parameter-Based DSMs
Unmanned Aerial Vehicle Microfilm Cartridge Heat Exchanger Automobile Brake System
1 Mixtral-7x8B with 12.5±2.2 19.5±5.0 16.1±2.2 12.1±3.1
without 15.8±4.6 17.7±3.3 16.8±1.9 16.0±2.3
Llama3-70B with 7.6±0.8 8.3±0.5 7.9±2.0 6.1±1.1
without 12.5±3.8 10.2±1.2 10.1±1.8 7.0±2.2
GPT-4-Turbo with 9.6±3.2 10.0±1.6 9.2±2.1 5.0±1.2
without 13.9±4.3 11.0±1.5 12.9±3.1 7.5±2.7
DeepSeek-V3 with 7.3±0.5 8.0±0.0 6.6±1.0 3.9±0.5
without 9.3±2.0 8.7±0.5 8.1±1.0 5.8±2.0
Claude-3.5-Sonnet with 6.6±0.7 8.0±0.0 4.8±0.6 4.4±0.9
without 11.9±3.4 8.1±0.3 5.8±1.1 5.9±1.4
5 Mixtral-7x8B with 10.6±1.4 13.1±2.7 13.0±2.1 9.7±2.5
without 12.3±2.6 15.0±2.2 15.4±2.3 13.8±1.7
Llama3-70B with 6.7±0.5 8.1±0.3 5.5±0.9 5.4±1.1
without 8.4±1.7 9.3±0.5 8.7±1.2 5.4±1.4
GPT-4-Turbo with 6.8±0.6 8.4±0.5 6.6±1.3 4.6±0.9
without 8.9±0.8 9.6±0.8 10.1±2.7 5.7±1.8
DeepSeek-V3 with 7.0±0.0 8.0±0.0 5.5±0.5 3.3±0.5
without 7.2±2.4 8.7±0.5 7.9±1.0 4.6±1.4
Claude-3.5-Sonnet with 6.1±0.3 8.0±0.0 4.0±0.4 3.0±0.0
without 7.5±0.9 8.0±0.0 4.9±0.5 4.0±1.0
20 Mixtral-7x8B with 10.1±1.2 11.2±2.3 10.7±1.2 8.8±2.0
without 11.3±1.7 13.0±1.2 14.0±1.3 13.1±1.7
Llama3-70B with 6.7±0.5 8.0±0.0 4.7±0.5 5.2±1.2
without 7.4±0.8 8.8±0.6 7.7±1.3 4.6±0.9
GPT-4-Turbo with 6.1±0.3 8.2±0.4 5.1±0.8 4.1±0.7
without 7.3±0.6 9.1±0.3 7.7±1.8 5.0±1.5
DeepSeek-V3 with 7.0±0.0 8.0±0.0 5.4±0.5 3.1±0.3
without 7.0±0.0 8.7±0.5 7.8±1.1 4.0±1.3
Claude-3.5-Sonnet with 6.0±0.0 8.0±0.0 3.6±0.5 3.0±0.0
without 6.4±0.7 8.0±0.0 4.1±0.3 3.4±0.7
5.4 Implications for Engineering Design
So far, the experimental results have demonstrated that our LLM-based combinatorial optimization method is effective and efficient across various DSM cases. To illustrate the reasoning process of our method, we visualize a representative trajectory from the Heat Exchanger DSM optimization process (Figure 4), where the LLM progressively improves the node sequence over four iterations. This visualization demonstrates how the model iteratively learns and refines solutions. In the first matrix randomly generated (Iteration 0), the initial node ordering yields 25 feedback loops, indicating a high level of interdependency and inefficiency. By the first iteration, the LLM suggests a new configuration that already reduces the feedback loop count to 5. The third and final iteration further optimizes the sequence, achieving the minimal known value of 4 feedback loops. With the optimization process, upstream nodes such as “Requirements” (N1), “Prognosis” (N3), and “Type of Concept” (N9) are consistently placed early, aligning with their role in initiating design decisions. Downstream elements like “Tool Manufacturing” (N17) and “Prototype Series” (N12) appear later in the sequence, preserving a logical and efficient flow of information. In practice, such visualization of the optimization trajectory enables us to analyze and trace the reasoning behind the LLM’s decisions, thereby enhancing the interpretability of the proposed method.
Figure 4: Optimization trajectory on the Heat Exchanger DSM case using the proposed method
It is important to note that for any given DSM, multiple optimal solutions may exist [1]. With the proposed method, users can either select from the top candidates in the final solution base or rerun the method with different seeds to generate diverse and high-quality alternatives. This flexibility is particularly valuable in engineering contexts, where different project goals or constraints may influence the selection of one optimal solution over others.
In sum, the proposed method can directly contribute to design planning and system-level coordination in engineering design practice. For example, in activity-based DSMs, it supports early-phase decision-making by suggesting task sequences that mitigate rework and streamline communication. In parameter-based DSMs, it reduces technical complexity by identifying sequencing patterns that minimize feedback loops. By leveraging both system structure information and domain knowledge, our method enables the LLM to reason in alignment with engineering logic, rather than relying solely on statistical correlations. This results in more effective and efficient optimization outcomes, ultimately accelerating and enhancing the overall engineering design process.
While this study focuses on sequencing problems within DSMs, the proposed approach can be generalized to a broader class of CO problems [57, 58, 59] in engineering and more fields. A more generalized framework is illustrated in Figure 5. During the engineering design process, systems or tasks can often be formulated as iterative optimization problems over formal representations, where multimodal information (such as task descriptions, parameter interactions, or architectural constraints encoded in text, diagrams, or code format) is encoded and integrated. The framework captures this pattern by transforming such information into a combined input information and integrating them with objective evaluation in a closed-loop reasoning cycle, enabling generation-based optimization using foundational AI models.
Figure 5: Foundational AI Model-based Framework for Solving General CO Problems
6 Limitations and Future Work
Although the proposed method exhibits strong performance, several limitations remain. First, the current evaluation is based on a relatively small number of DSM cases. Expanding the dataset to include larger and more complex systems from diverse engineering domains would allow for more comprehensive analysis and improved statistical reliability. Second, this study focuses specifically on sequencing problems, whereas the proposed framework may be applicable to a wider range of CO problems in engineering design, such as modularization and change propagation analysis [57, 58, 59]. As presented in Figure 5, the framework is designed to generalize beyond sequencing by incorporating structured system information and domain-specific semantics into an iterative optimization process. Future work may explore the effectiveness and adaptability of this approach across these expanded tasks.
Another important direction for future research lies in improving the interpretability of the optimization process. In this paper, we presented a preliminary case illustrating how the LLM refines the DSM through a sequence of iterative updates. However, the underlying decision-making and reasoning mechanisms of LLMs still remain relatively opaque. Researchers may explore more effective techniques for LLM interpretability, such as integrating complementary models, enabling self-explanation, and visualizing decision pathways [60].
Additionally, the current framework primarily relies on textual representations of domain knowledge and system structures. While effective, this may limit its ability to fully leverage non-textual modalities such as diagrams, tables, or structured data formats commonly used in engineering practice. In future work, we will explore Multimodal Large Language Models (MLLMs) [61] to incorporate visual or tabular information, enriching the reasoning input and expanding the framework’s applicability to more complex engineering workflows.
7 Conclusion
In this paper, we have proposed a knowledge-informed optimization approach that leverages LLMs to perform iterative combinatorial optimization on DSMs. By integrating structural network information with contextual domain knowledge, the proposed method leverages the advanced reasoning capabilities of LLMs to iteratively generate and refine high-quality solutions. Experiments conducted on various DSM cases demonstrate that the proposed approach outperforms both stochastic and deterministic baselines in terms of convergence speed and solution quality. The consistent performance gains observed when incorporating domain knowledge further reveal the value of combining semantic insight with mathematical reasoning. These results suggest that LLMs provide a promising foundation for a new paradigm in design optimization and decision support, particularly within the complex and knowledge-intensive environments of engineering design.
Appendix 1. Full prompts
Prompt for the proposed approach:
You are an expert in the domain of combinational optimization. Please assist me to find an optimal sequential order that minimizes feedback cycles in the dependency network described below. Your task is to propose a new order that differs from previous attempts and has fewer feedback cycles than any listed. <Description of the Entire Network> {network_description} </Description of the Entire Network>
<Nodes with Descriptions>
{node_list_with_description} </Nodes with Descriptions>
<Edges>
{edge_list} </Edges>
Below are some previous sequential orders arranged in descending order of feedback cycles (lower is better): {selected_historical_solutions} Please suggest a new order that: - Is different from all prior orders. - Has fewer feedback cycles than any previous order. - Covers all nodes exactly once. - Starts with <order> and ends with </order>. - You can use the descriptions of nodes and networks to support your suggestion. Output Format: <order> ...... </order> Please provide only the order and nothing else.
Prompt for the proposed approach (removing knowledge):
You are an expert in the domain of combinational optimization. Please assist me to find an optimal sequential order that minimizes feedback cycles in the dependency network described below. Your task is to propose a new order that differs from previous attempts and has fewer feedback cycles than any listed. <Nodes> {node_list} </Nodes>
<Edges>
{edge_list} </Edges>
Below are some previous sequential orders arranged in descending order of feedback cycles (lower is better): {selected_historical_solutions} Please suggest a new order that: - Is different from all prior orders. - Has fewer feedback cycles than any previous order. - Covers all nodes exactly once. - Starts with <order> and ends with </order>. Output Format: <order> ...... </order> Please provide only the order and nothing else.
Example of {network_description} in the UCAV design activity DSM case:
This network represents the dependency relationships among conceptual design activities for UCAV development at Boeing. Each node corresponds to a specific task or analysis, and directed edges indicate the prerequisite relationships between these tasks. Nodes: Each node is a task or analysis in the conceptual design process. Edges: Directed edges show the prerequisite relationships between these tasks and analyses.
Example of {node_list_with_description} in the UCAV design activity DSM case:
{’id’: ’lzOtR’, ’name’: ’Create Configuration Concepts’},
{’id’: ’yLlKi’, ’name’: ’Prepare UCAV Conceptual DR&O’},
{’id’: ’Swvi2’, ’name’: ’Prepare 3-View Drawing & Geometry Data’},
{’id’: ’CDcxF’, ’name’: ’Perform Weights Analyses & Evaluation’},
{’id’: ’0KGDm’, ’name’: ’Perform Aerodynamics Analyses & Evaluation’},
{’id’: ’4wHtv’, ’name’: ’Perform Multidisciplinary Analyses & Evaluation’},
{’id’: ’AgIBP’, ’name’: ’Prepare & Distribute Choice Config. Data Set’},
{’id’: ’gRtHi’, ’name’: ’Perform S&C Characteristics Analyses & Eval.’},
{’id’: ’GV9RJ’, ’name’: ’Make Concept Assessment and Variant Decisions’},
{’id’: ’I1j2m’, ’name’: ’Perform Performance Analyses & Evaluation’},
{’id’: ’Vzzm7’, ’name’: ’Perform Propulsion Analyses & Evaluation’},
{’id’: ’B0BFG’, ’name’: ’Perform Mechanical & Electrical Analyses & Eval.’}
Example of {node_list} in the UCAV design activity DSM case:
[’lzOtR’, ’yLlKi’, ’Swvi2’, ’CDcxF’, ’0KGDm’, ’4wHtv’, ’AgIBP’, ’gRtHi’, ’GV9RJ’, ’I1j2m’, ’Vzzm7’, ’B0BFG’]
Example of {edge_list} in the UCAV design activity DSM case:
{’dependent’: ’0KGDm’, ’predecessor’: ’Swvi2’},
{’dependent’: ’AgIBP’, ’predecessor’: ’lzOtR’},
{’dependent’: ’0KGDm’, ’predecessor’: ’yLlKi’},
{’dependent’: ’Swvi2’, ’predecessor’: ’lzOtR’},
Example of {selected_historical_solutions} in the UCAV design activity DSM case:
{’solution’: ’lzOtR, yLlKi, GV9RJ, AgIBP, B0BFG, Vzzm7, Swvi2, CDcxF, 0KGDm, I1j2m, gRtHi, 4wHtv’, ’score’: 15.0},
{’solution’: ’B0BFG, yLlKi, Vzzm7, lzOtR, Swvi2, CDcxF, AgIBP, 0KGDm, GV9RJ, I1j2m, gRtHi, 4wHtv’, ’score’: 13.0},
Appendix 2. Parameter settings for three variants of GA
In all three variants of the GA used for DSM sequencing tasks, we employed the following shared settings: the number of generations was set to 2,000, with a selection mechanism using tournament selection and mutation using shuffled indexes. The GA was implemented using the DEAP library with standard configurations for initial population generation. The different configurations for each GA setting are shown in the table below. The meaning of each parameter also refers to [49].
Population Individual Mutation Probability Tournament Size Crossover Probability Mutation Probability
Exploration-focused 50 0.05 5 0.6 0.4
Exploitation-focused 10 0.01 20 0.9 0.1
Balanced 20 0.02 10 0.7 0.3